1.) A race car accelerates uniformly from 18.5 m/s to 46.1 m/s in 2.47 seconds. Determine the acceleration of the car and the distance traveled.

2.) A feather is dropped on the moon from a height of 1.40 meters. The acceleration 

of gravity on the moon is 1.67 m/s

2. Determine the time for the feather to fall to the surface of the moon.

3.) An engineer is designing the runway for an airport. Of the planes that will use the 

airport, the lowest acceleration rate is likely to be 3 m/s². The takeoff speed for this plane will be 65 m/s. Assuming this minimum acceleration, what is the minimum allowed length for the runway?


1
Expert's answer
2022-03-26T15:31:03-0400

1.

"a=\frac{v_2-v_1}t=11.2~\frac m{s^2},"

"s=\frac{v_2^2-v_1^2}{2a}=79.8~\frac ms."

2.

"t=\sqrt{\frac{2s}g}=1.3~s."

3.

"S=\frac{v^2}{2a}=700~m."